Jane is Y years old on the 31st of January 2015. Jane's mother is 24 years older than her. If Jane will be half her mother's age in 10 years time, in what year was Jane born?
Answer by lwsshak3(11628)

Jane is Y years old on the 31st of January 2015. Jane's mother is 24 years older than her. If Jane will be half her mother's age in 10 years time, in what year was Jane born?
***
31 Jan 2015:
y=Jane's age
y+24=Her mother's age
in 10 years:
y+10=Jane's age
y+24+10=y+34=her mother's age
y+34=2(y+10)
y+34=2y+20
y=14
in what year was Jane born? 2001